OpenStudy (anonymous):

The equation, y = 10(1.0605)x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000.

OpenStudy (campbell_st):

is the equation \[y = 10(1.0605)^x\]

OpenStudy (anonymous):

Yeahh lol I forgot the ^, sorry :p

OpenStudy (campbell_st):

ok... so in 1911 its \[y = 10 \times (1.0605)^0\] when you get a numerical answer multiply it by 1000 since the information says... in thousands of dollars. so really you could write the equation for 1911 as \[y = 10 \times (1.0605)^0 \times 1000\] the just evaluate

OpenStudy (anonymous):

Ohh ok thank you so much :)